OpenStudy (anonymous):

Malorie can work 10 math problems in 3 minutes. At this rate, how many problems can she work in one hour? A. 30 B. 60 C. 120 D. 200

OpenStudy (thecatman):

60/3=20 ------ 20*10 =

OpenStudy (anonymous):

20*10= 200

OpenStudy (anonymous):

thanks @TheCatMan